OpenStudy (anonymous):

I need help in solving this question plz. Differentiate the following y=x (x^2+1)^4

OpenStudy (anonymous):

expand the equation ..

OpenStudy (anonymous):

no.. @pengembara_bumi1

OpenStudy (anonymous):

:( what's your answer @itsmeeQUENNIE

OpenStudy (anonymous):

@Ash_Hlovate

OpenStudy (anonymous):

y= x (x^8 + 4x^6 + 6x^4 + 4x^2 + 1)

OpenStudy (anonymous):

3x^2+1

OpenStudy (anonymous):

dy/dx = u(dv/dx) + v(du/dx) = x(2x)+(x^2+1)(1) = 3x^2+1

OpenStudy (anonymous):

re you sure with that @Ash_Hlovate ?

OpenStudy (anonymous):

haha.. i think so. but if im wrong.. im sorry. ;)

OpenStudy (anonymous):

great work @Ash_Hlovate ! :D

OpenStudy (anonymous):

(x^2 +1)^3 . (8 x^3 + x^2 +8x+1). Is it correct.